A ball has a mass of 2kg.It is dropped from a cliff and strikes the ground below at 10m/s.


1) Which is the kinetic energy as it is about to strike the ground?Ek=mv2


2) What was its potential energy before it was dropped? Ep=mgh


3) Determine the height from which was it was dropped? E=mgh



1
Expert's answer
2021-05-28T07:16:08-0400

Solution.

"m=2kg;"

"v=10m\/s;"

"1) E_k=\dfrac{mv^2}{2};"

"E_k=\dfrac{2kg\sdot100m^2/s^2}{2}=100J;"

"2)E_p=E_k=100J;"

"3)E_p=mgh\implies h=\dfrac{E_p}{mg};"

"h=\dfrac{100J}{2kg\sdot9.8m/s^2}=5.1m;"

Answer:"1)100J;"

"2)100J;"

"3)5.1m."